Calculate the volume of hydrogen sulphide measured at 15 degree celcus and 1 atm pressure which will react with a solution containing 4g of copper sulphide


1
Expert's answer
2021-06-14T13:35:25-0400

PV = n RT

n = 4/95.611 = 0.041 mol

T = 15⁰C = 288K

1×V = 0.041 × 0.0821 × 288

V = 0.96L